[Hallc_running] [New Logentry] RC update (Sept 12) -- NO RC meeting today

brads at jlab.org brads at jlab.org
Sat Sep 12 16:05:02 EDT 2020

Logentry Text:
<p>Dave G. and I will be taking an opportunistic access at 4:15pm today to try to address the iochc10 issue.</p><p><strong>We will cancel the 4:15pm meeting today. </strong> There will be a meeting tomorrow (Sunday)</p><p><strong>NOTE: I was recently notified by the PD that the Hall C controlled access on Monday may shift to 2pm instead of 10am. </strong></p><hr><p><strong>Saturday 12 Sept 2020. </strong></p><h1><span id="General_updates" class="mw-headline">General updates</span></h1><p>(See <a title="D2n Run Plan 2020" href="https://hallcweb.jlab.org/wiki/index.php/D2n_Run_Plan_2020">d2n Run Plan 2020</a> for broader status/overview.)</p><ul><li>Presently down from 2pm (until 5--6pm) for an RF klystron replacement [https://logbooks.jlab.org/entry/3846701]</li><li>Production running on Transverse target (@ 90); Kin-C, Kin-Z continues.<ul><li>General plan when this is done is to complete Kin-A, Kin-X (partially done already); then move on to Kin-B/Y and/or 1-pass r!
 unning as time allows.</li></ul></li><li><strong>Moller</strong> measurement was rescheduled for today (2nd attempt); still working on it as of 3:30pm (Accelerator issues) <a class="external autonumber" href="https://logbooks.jlab.org/entry/3846082" rel="nofollow">[1]</a></li><li>Should move to Kin-A, Kin-X in the next day or so and finish 90-deg running there</li><li>Target should be rotated to 270 deg on Monday</li><li>Max Target Tommy polarization @ 90 since yesterday has been 38.5% <a class="external autonumber" href="https://logbooks.jlab.org/entry/3845852" rel="nofollow">[2]</a>, it seems to have plateaued at 378% today <a class="external autonumber" href="https://logbooks.jlab.org/entry/3846692" rel="nofollow">[3]</a>, <a class="external autonumber" href="https://logbooks.jlab.org/entry/3846581" rel="nofollow">[4]</a><ul><li>For comparison, maximum seen with target at 90 deg last week was 44% (9/5/2020 22:56:38)</li><li>FWIW, that is similar to the maximum seen at 27!
 0 was 37.5% <a class="external autonumber" href="https://logbo!
 oks.jlab..org/entry/3844366" rel="nofollow">[5]</a>; Recent max around 36% (Thursday) <a class="external autonumber" href="https://logbooks.jlab.org/entry/3845243" rel="nofollow">[6]</a></li></ul></li><li>Moller measurement completed Friday 86.1% <a class="external autonumber" href="https://logbooks.jlab.org/entry/3846252" rel="nofollow">[7]</a>, (incl. 0.9% bleedthrough correction from Hall A)</li><li>Significant down during Friday swing due to problem with iochc10 IOC which handles Moller controls, magnet cryo, etc<ul><li>Found to be in boot, crash, reboot loop <a class="external autonumber" href="https://logbooks.jlab.org/entry/3846242" rel="nofollow">[8]</a>, hardware and software investigations (including Software On-Call) ultimlately unsuccessful <a class="external autonumber" href="https://logbooks.jlab.org/entry/3846365" rel="nofollow">[9]</a>. System went down again in Owl <a class="external autonumber" href="https://logbooks.jlab.org/entry/3846525" rel="nofollow">!
 [10]</a></li><li>Next step is likely swapping out the VME crate; talking to Jack for some support.</li><li>Brad will install some cameras to at least monitor the cryo situation at the next opportunity</li></ul></li><li>Once the Hall came out of Controlled Access in early swing beam has been pretty good</li></ul><ul><li>Junhao reports the spectrum analyzer parts are ready to go for Monday<ul><li>RadCon has been looped in and will support</li></ul></li></ul><h2><span id="Notes_for_the_Future" class="mw-headline">Notes for the Future</span></h2><ul><li>Beam Studies planned for <strong>Monday</strong>@ 10am to 6pm (including recovery), Sept 15<ul><li>RadCon will support going to CA at 10am, we can stay in CA for roughly 2 hours; OPS wants Hall C dump available after that</li><li>Simona wants access at this time for SOLID cherenkov related work</li><li>Target work<ul><li>Install spectrum analyzer collimator</li><li>Check field at target region for 270 setting, if possible</li><l!
 i>Want to determine field direction at the precision of a few degrees vs!
  the alignment laser</li></ul></li><li>Bill, Murchhana check compass/field measurements and will report back at Sat RC meeting</li></ul></li></ul><ul><li><a title="D2n Run Plan 2020" href="https://hallcweb.jlab.org/wiki/index.php/D2n_Run_Plan_2020#1-pass_Running">1-pass running</a>tentatively scheduled for Sept 17/18<ul><li>Hall A would like a specific date soon so they can finalize their end-of-run plans</li></ul></li></ul><h2><span id="Pending_Issues" class="mw-headline">Pending Issues</span></h2><ul><li>Check EPR calibrations (particularly for 270deg setting, but others too) [Melanie]</li><li>Magnet EPICS PV timeout issue; may have to reboot skylla10 on Monday?<ul><li>Some Magnets Channels Lost Connection Briefly <a class="external autonumber" href="https://logbooks.jlab.org/entry/3846677" rel="nofollow">[11]</a> (Saturday, 2pm)</li><li>Also had to add more error checking to the angle reporting in the CODA run-start GUI last week too (same softIOC as above)</li><li>'Hear!
 tbeat' issue reported by Steve Lassiter last week</li></ul></li><li>Target polarization for Tommy @ 270 is low (mid to high 30s)<ul><li>Spectrum analyzer collimator was broken at some point and present optics setup may be allowing transient light to interfere with spectrum measurement and QWP scans</li><li>JP wonders if the field direction is off for the transverse settings (Bill, Murchhana: see note in Target group chat from late Thursday)</li></ul></li></ul><ul><li>Beam energy drifts<ul><li>Hall C 5-pass polarization goes down as energy (HALLC:p) goes up <a class="external autonumber" href="https://logbooks.jlab.org/entry/3844493" rel="nofollow">[12]</a></li></ul></li></ul><h2><span id="Opportunistic_Jobs.2FAccess" class="mw-headline">Opportunistic Jobs/Access</span></h2><ul><li>Beam Studies planned for <strong>Monday</strong>, Sept 14 (Note date change!)<dl><dt>Target calibrations / tests</dt></dl><ul><li>Just before Moller measurement, and/or other known downtimes</li><!
 li>See <a title="D2n Run Plan 2020" href="https://hallcweb.jlab.org/wik!
 i/index.php/D2n_Run_Plan_2020#Target_Calibrations_and_Tests">D2n_Run_Plan_2020#Target_Calibrations_and_Tests</a></li></ul></li></ul><h1><span id="Current_Run_Plan" class="mw-headline">Current Run Plan</span></h1><h2><span id="Saturday_DAY_.E2.86.92_Saturday_SWING" class="mw-headline">Saturday DAY → Saturday SWING</span></h2><ul><li><strong>Continue</strong> production data (HMS Kin-C/SHMS Kin-Z) with pol 3He target, <strong>Transverse/90</strong> polarization with NMR every 5 hrs.</li><li>[TBD] Move to Kin-A, Kin-X and continue Transverse/90 running there?</li></ul><h2><span id="Saturday_OWL_.E2.86.92_Sunday_DAY" class="mw-headline">Saturday OWL → Sunday DAY</span></h2><ul><li>Continue production data with pol 3He target, <strong>Transverse/90</strong> polarization with NMR every 5 hrs.</li></ul>


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3846729
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20200912/565456cc/attachment.html>

More information about the Hallc_running mailing list